Reduce 2++x²+2x₂x4x−2x4, into canonical form. Find the rank,





index, signature and its nature.


1
Expert's answer
2021-12-09T03:08:58-0500

The reduction of a quadratic can be carried out by a procedure known as Lagrange’s Reduction, which consists essentially of repeated completing of the square:

"q=x_1^2+2x^2_2-7x_3^2-4x_1x_2+8x_1x_3=(x_1^2-4x_1(x_2-2x_3))+2x_2^2-7x_3^2="

"=(x_1^2-4x_1(x_2-2x_3)+4(x_2-2x_3)^2)+2x_2^2-7x_3^2-4(x_2-2x_3)^2="

"=(x_1-2x_2+4x_3)^2-2(x_2^2-8x_2x_3)-23x_3^2="

"=(x_1-2x_2+4x_3)^2-2(x_2^2-8x_2x_3+16x_3^2)+9x_3^2="

"=(x_1-2x_2+4x_3)^2-2(x_2-4x_3)^2+9x_3^2"


"y_1=x_1-2x_2+4x_3"

"y_2=x_2-4x_3"

"y_3=x_3"


"q=y_1^2-2y_2^2+9y_3^2"


index is  the number of positive terms in the canonical form:

"p=2"

the signature is number of positive terms minus the number of negative terms:

"n=2-1=1"

rank:

"r=n+p=3"

nature:

there is a mixture of positive and negative terms in the canonical form, so this is indefinite form